Speedy Auto Repairs uses a job-order costing system. The company’s direct materials consist of replacement parts installed in customer vehicles, and its direct labor consists of the mechanics’ hourly wages. Speedy’s overhead costs include various items, such as the shop manager’s salary, depreciation of equipment, utilities, insurance, and magazine subscriptions and refreshments for the waiting room.

The company applies all of its overhead costs to jobs based on direct labor-hours. At the beginning of the year, it made the following estimates:

Direct labor-hours required to support estimated output 16,000
Fixed overhead cost $ 168,000
Variable overhead cost per direct labor-hour $ 1.00

Required:

1. Compute the predetermined overhead rate.

2. During the year, Mr. Wilkes brought in his vehicle to replace his brakes, spark plugs, and tires. The following information was available with respect to his job:

Direct materials $ 622
Direct labor cost $ 229
Direct labor-hours used 6

Compute Mr. Wilkes’ total job cost.

3. If Speedy establishes its selling prices using a markup percentage of 60% of its total job cost, then how much would it have charged Mr. Wilkes?
Answer #1

Total variable overhead estimated=(16000*1)=$16000

Total overhead estimated=Total variable overhead estimated+Total fixed overhead estimated

=16000+168,000=$184,000

a.Predetermined overhead rate=Total overhead estimated/Estimated direct labor hours

=184,000/16000

=$11.5 per labor hour

b.Total overhead applied=Predetermined overhead rate*Actual direct labor hours

=11.5*6=$69

Total job cost=Direct materials+Direct labor+Total overhead

=622+229+69

=$920

c.Selling price=Total job cost*160%

=920*160%

=$1472